OpenStudy (anonymous):

How do I graph a circle with this equation: (x^2)+(y^2)-2x+4y-20=0

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

Sort of

OpenStudy (anonymous):

I know what it is but I don't know how to turn it into that

OpenStudy (anonymous):

Please help

OpenStudy (jdoe0001):

well, let's start by grouping \(\bf x^2+y^2-2x+4y-20=0 \implies (x^2-2x)+(y^2+4y)-20=0\quad \\\quad \\ (x^2-2x+\square?)+(y^2+4y+\square?)-20=0\) what values do you think we need in each group, to get a perfect square trinomial? what's the missing value in each?

OpenStudy (anonymous):

-8 and -12 maybe? I started grouping already

OpenStudy (anonymous):

But I included the -20 in mine

OpenStudy (jdoe0001):

hmm right.. so... \(\bf (x^2-2(\square)x+\square^2)+(y^2+2(\square)y+\square^2)-20=0\) what do you think we need now?

OpenStudy (anonymous):

4 and 6

OpenStudy (anonymous):

Does that sound right?

OpenStudy (jdoe0001):

out middle term in a perfect square trinomial is the product of the other 2 terms, without the square

OpenStudy (jdoe0001):

so, can't be 4 or 6, in either case

OpenStudy (jdoe0001):

product of the other 2 terms, without the square * 2 , that is

OpenStudy (anonymous):

But that can't be

OpenStudy (anonymous):

Could you show me what you would do and explain how you got it?

OpenStudy (anonymous):

I am getting more confused

OpenStudy (jdoe0001):

lemme do the 1st one... \(\bf (x^2-2x+\square^2)\\ 2 \times x \times \square = 2x\\ \square = 1\\ (x^2-2x+\square^2) \implies (x^2-2x+1^2) \implies (x-1)^2\)

OpenStudy (anonymous):

Is the second one (y+1)^2?

OpenStudy (jdoe0001):

\(\bf \large{ (y^2+4y+\square^2)\\ 2 \times y \times \square = 4y}\)

OpenStudy (jdoe0001):

well... not quite 1

OpenStudy (anonymous):

2

OpenStudy (anonymous):

(y-2)^2

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

\(\bf (y^2+4y+\square^2)\\ 2 \times y \times \square = 4y\\ \square = 2\\ (y^2+4y+\square^2)\implies (y^2+4y+2^2) \implies (y+2)^2\)

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

So it should be (x-1)^2 + (y+2)^2 = 20?

OpenStudy (jdoe0001):

notice though, we added 1 first, for one, and then 4 for the other what we are doing is just borrowing from zero, so if we ADD 1 and 4, we have to also subtract 1 and 4 so \(\bf x^2+y^2-2x+4y-20=0 \implies (x^2-2x)+(y^2+4y)-20=0\quad \\\quad \\ (x^2-2x+1^2)+(y^2+4y+2^2) - 1^2-2^2-20=0\\\quad \\ (x-1)^2+(y+2)^2- 25=0 \implies(x-1)^2+(y+2)^2 = 25\)